# Predictive Pattern Recognition Techniques Towards Spatiotemporal Representation of Plant Growth in Simulated and Controlled Environments: A Comprehensive Review

### v2

@@ 4.1 Functional-Structural Plant Models
Functional-structural plant models (FSPM) represent a class of plant representation approaches to model the structural and geometric elements of a plant through its mechanistic processes [Godin2005, 10.1093/jxb/erp345]. Particularly, these models capture the interaction between the plant’s physiological processes, such as photosynthesis, water transport, and nutrient allocation, and its development of plant organs such as leaves, branches, and roots over time [10.1093/aob/mcx144, Evers2016, Xiujuan2024, doi:10.1073/pnas.1410238111]. By coupling these two aspects, FSPMs provide a comprehensive view of plant behavior in response to various dependent parameters, allowing for the visualization of resource allocation and its effect on growth patterns. More recently, empirical methods have been integrated to the development process of FSPMs to establish the relationships between features extracted from HTPP data [doi:10.34133/plantphenomics.0127]. As such, an emergence of hybrid data-driven, mechanistic FSPMs have been employed for plant growth simulations.
@@ 4.1.1 Process-driven Differential Growth Models
One application of these FSPMs is to capture the morphological developments within plant organs through differential growth models. The development and emergence of complex 3D morphologies within plant structures through time can be represented through numerical models, such as finite element methods (FEM), to describe differential growth across localized regions. The basis of the mechanistic models that determine these differentials are typically derived through experimental observations, utilizing image-processing techniques such as optical flow to obtain growth parameters from HTPP data [10.3389/fpls.2019.00227]. Changes are mapped spatially and reconstructed from the distribution of conditions which influence the final shape of plant organs. One example of such modeling is demonstrated in Huang et al. [huang2018differential], where biomechanical conditions are observed to determine the shape formations of leaves and petals. In this work, the modeling of growth differentials can be used to describe and reconstruct twisting, helical twisting, and edge-waving configurations within leaf structures through FEM simulations.
More often, this approach for plant structural modeling is applied toward developing fate and growth maps of cells within individual plant organs to simulate localized development [Mosca2018]. For instance, in the work done by Kierzkowski et al. [KIERZKOWSKI20191405], the dynamics of genomic-proteomics pathways of two homeobox genes and their effects on promoting, inhibiting, and differentiating plant cells are modeled through spatial growth differentials using FEM. This phenomena is modeled through time and driven by the localized stimuli of plant hormone, namely auxin, concentrations. In this case, this model is applied toward developing a deep understanding of the causal effects of these specific signal pathways through empirical methods by mapping signals through time.
4.1.2 Rule-based Lindemyer-Systems
Another common implementation of FSPMs involves the application of a rule-based modeling paradigm known as the Lindemyer-systems (L-systems). L-systems define a formal set of recursive rules that specify the configuration of plant structures based on a string of symbolic characters. These rules guide the growth and branching patterns of simulated plants through iterative processes that aim to generalize and represent the realistic biological development and propagation of plant structures [prusinkiewicz2012algorithmic, Prusinkiewicz2018]. By repeatedly applying these recursive rules, L-systems can simulate natural variations and intricate details found in plant morphology, making them a powerful tool in computational botany and generative modeling.
A basic example is given by Algorithm 1, which establishes a generalized execution of a simple L-system, where Figure 6 shows the visual rendering of a rule set commonly used for describing a simple branching pattern in 2D.
@@ 4.1.2 Rule-based Lindemyer-Systems
This algorithm can be expanded to include parameterized production rules that encapsulate complex relationships at each iteration of the simulated plant branching. For example, various works have developed production rules that dictate tree organ formation based on a source-sink dynamic between existing plant organs at previous iterations, which is then subsequently rendered for the 3D simulation of various tree cultivars [LPeaches2005, dejong2015almond]. More recent work has expanded on production rules using sink-source dynamics by incorporating context-driven systems such as environmental and management conditions, soil nutrient availability, and distribution to improve the dynamic tree simulation process [LPeaches2022, 9231935, YANG2024122].
Furthermore, works such as Hitti et al. [HITTI2024108613] and Mußmann et al. [10.1162/isal_a_00718] introduce an adaptation of the L-systems paradigm that captures the dynamic development of plant structures using agent-based interactions. Plant structures, represented as agents, can independently interact with various stimuli. Within such systems, agent-based feedback mechanisms, such as reinforcement learning, can be used to train model growth patterns based on external stimuli.
@@ 4.2 Computer Vision Approach to Plant Modeling
FSPMs typically rely on procedural methods for the visual rendering process of the plant simulation pipeline (Figure. 5). Although the outcomes of such models are interpretable and result in visually consistent structures, the outputs of such methods are often derived from deterministic solutions that lack stochastic variability. Such solutions tend to drift from biological realism, particularly when simulating complex interactions or environmental responses, requiring the integration of experiment data to drive realistic outputs [10.1093/jxb/erp345]. Alternatively, some initial efforts have been made towards a novel approach using data-driven CV for rendering the structural elements of the simulation task.
@@ 4.2.1 Next Frame Prediction
Some early works on CV-based plant growth forecasting include Sakurai et al. [sakurai2019plant] and Wang et al. [agronomy12092213], which have proposed the adoption of convolutional recurrent neural networks (CRNN) for the task of image-to-image translation to predict the next frame of a plant growth scene given a set of prior frames from the same scene. This approach expands on the sequence-based prediction techniques described in Section 3.2.2 to assess spatial features from consecutive images. This is done by incorporating convolutional layers within a RNN framework to extract and encode features from a set of input images, $X$, to a transitional hidden state $h$. The rendering of this hidden state is described by a decoding step which consists of an up-sampling process of $h$ through deconvolutional layers. The typical generic implementation of the CRNN is shown in Figure 7. This approach combines pattern recognition techniques for forecasting plant parameters using RNNs with the visual rendering of plant structures through image translation methods. While CRNNs allow for spatial features to be captured through time, the deterministic mapping through convolutional and deconvolutional layers often result in artifacts and ambiguities within structures of the output images [kingma2013auto].
To mitigate problems in image quality, recent work have focused on improving the image generation component of model architectures, which correspond to the decoder of a CRNN. Yasrab et al. [rs13030331] proposes the use of a Generative Adversarial Network (GAN)-based network architectures for improved realistic image synthesis. Other works such as Kim et al. [9750078] propose multi-step networks that separate various components of the image generation task, such as shape estimation and image reconstruction, utilizing a transformer-based network for spatial pattern recognition. Although these approaches improve image generation quality, they do not address the challenges of ambiguous structures caused by a deterministic mappings. CV-based plant forecasting is an ill-posed problem, in which, there exists non-unique outcomes of a plants growth trajectory given the same initial set of frames. This is particularly evident under diverging environmental conditions. Moreover, predictions from the current models are made solely on the previous frames of the scene and further contribute to the prevalence of ambiguities during image translation.
@@ 4.2.2 Frame Synthesis Through Conditional Generative Models
A probabilistic approach can be taken to regularize the plant growth modeling problem and to guide the prediction process by incorporating stochastic elements caused by dynamic growing conditions. In such methods, a generative model is employed for the task of frame synthesis to predicted the expected frames of the plant growth scene. While generative models are typically used as data augmentation techniques for enhancing sparse agricultural datasets [doi:10.34133/plantphenomics.0041, LU2022107208], few works have explored the use of conditional generative models (CGM) for guided image synthesis. A probabilistic approach using a CGM is illustrated by a generic conditional variational autoencoder (CVAE) shown in figure 8. Instead of performing direct deterministic mappings between inputs and output frames, the objective of the CVAE is to learn a prior distribution, $p_{\theta}(z|x)$, where $z\sim p_{\theta}(z|x)$ is the latent representation of scene at a given point in time, $t$, directed by some conditional vector $x$. The prior distribution can be sampled such that it incorporates stochasticity through auxiliary variable, $\epsilon$. The network structure explored within the domain of plant growth modeling typically consists of a probabilistic latent variable encoder and a deterministic decoder to render output frames, $\hat{y}\sim p_{\theta}(y|z,x)$.
Early works, including Miranda et al. [9956115] and Drees et al. [DREES2021106415], propose the adoption of a conditional GAN (cGAN), which introduces a discriminator with adversarial routines to the generic CVAE structure, for synthesizing field images and forecasting the state of the crop canopy. The conditional vector consists of fixed soil treatment embeddings set at the start of the growing season. Forecasts are made at discrete points in time driven by the initial states of the treatment conditions. These works have demonstrated SOTA guided image generation for realistic image synthesis with variability in outputs defined by stochastic elements, or “factors of variation” (FOV). The frames synthesized from such models can be used in conjunction with image-based feature extraction models to provide quantitative estimation of biomass for a complete simulation process [10.1007/978-3-031-16788-1_29].
In these works, predictions are guided by a conditional vector comprised of treatment conditions, a prior frame state, and a time factor. The treatment embeddings are fixed at a particular point in time, typically set at the start of experimentation, and do not account for dynamic changes in conditions. For spatiotemporal simulation and trajectory estimation, the guidance of predictions without feedback can lead to larger variability when diverging from initial conditions due to the accumulated effects of compounding factors, including environmental conditions [Drees_2024]. As such, this approach may be applicable for simulating images of end of harvest canopies under varying treatment conditions at a discrete point in time in the future, but lacks the dynamics for real-time continuous plant growth predictions useful for real world applications such as anomaly detection and intervention practices.
Debbagh et al. [Debbagh2024SIPGS] proposes a sequence-informed approach which introduces dynamic and compounding environmental conditions as feedback to a CGM, guiding a continuous simulation process. In this work, the conditional vector, $x$, which guides the image generation of the CGM, is directed by a configuration of random variables, $c$, that describe dynamic factors that contribute to the development of the plants within the scene (i.e. environmental parameters). $x$ is determined by encoding a sequence of these environmental parameters from the start of the simulation up to the point of inference, $c(\tau)$, where $\tau=[0,t]$. The encoding process consists of an RNN encoder, which allows the model to learn the compounding and sequential relationships between environmental conditions as a vector representation. The dependencies between each model component is illustrated by the directed graph model shown in Figure 9. This approach provides a framework that further regularizes the image generation task through dynamic context, allowing for the realistic synthesis of consecutive frames.
@@ 5.1 Data-Driven vs Experiment-Driven Approaches
Plant growth modeling has traditionally relied on experiment-driven approaches, where predefined hypotheses are tested using carefully controlled experiments. These methods provide valuable insights into specific biological processes but are often time-consuming, resource-intensive, and limited in scalability. By contrast, data-driven and model-centric approaches leverage large datasets and machine learning techniques to capture patterns and relationships that might not be immediately apparent within an experiment-driven framework [roscher2023datacentricdigitalagricultureperspective]. The works presented in this review demonstrate a shift towards the latter, as challenges with dynamic updating and long-term predictions are prevalent in simple regression methods as described in Section 3.1.2.
A shift towards 2D image and 3D point cloud data for describing plant phenemena, as with HTPP datasets, requires model frameworks designed for interpreting and processing higher order data-structures. This inherently moves prospective research towards the CV-based models described in Sections 4.2.1 and 4.2.2. However, the advantages of these data-driven approaches come with the limitation of reduced interpretability and a lack of the biological grounding inherent in experiment-driven methods. In contrast, FSPMs provide a robust simulation framework for structured representation that is highly controllable, but is completely reliant on experiment-driven approaches for developing knowledge-based rule sets.
Future research in this domain should prioritize the development of hybrid approaches that integrate the strengths of both paradigms. Two key directions are proposed to achieve this goal. The first involves designing model architectures that incorporate knowledge-informed mechanisms, enabling the integration of biological principles and domain expertise into data-driven frameworks. The second focuses on advancing methods for decoding and interpreting opaque models, ensuring that complex representation models provide outputs that align with biological processes and facilitate deeper understanding.
@@ 5.2 Dataset Gap
The shift toward data-driven approaches has placed datasets at the center of plant growth modeling and predictive frameworks. Large-scale, high-quality datasets are necessary for training DL models and validating their predictions. However, the effectiveness of these models are dependent on the availability, diversity, and representativeness of the data. Despite advancements in HTPP platforms, significant gaps remain in the scope, quality, and accessibility of existing  datasets.
First, many datasets focus on single time points or static snapshots, which fail to capture the dynamic growth trajectories of plants over time. Existing publicly available datasets that capture longitudinal data (Table 1) are often scarce, and limited in sample frequency. This is particularly evident for datasets that capture multiple modalities of data including 2D images and 3D point clouds along with environmental parameters. Second, the diversity of crop species represented in current datasets is limited, with a heavy emphasis on staple crops such as maize (Zea mays) and soybean (Glycine max). This narrow scope restricts the ability of models to generalize across plant species with differing morphologies, such as root vegetables or perennials. Finally, the lack of consolidation of datasets across different research groups and platforms creates significant barriers to broader adoption and cross-comparative studies. Inconsistent formats, annotation practices, and metadata standards hinder the ability to combine datasets, thereby limiting the diversity and scale necessary for developing generalizable models.
To address these gaps, several considerations for developing future datasets are proposed. First, works should aim to apply standardization and open-access practices during dataset development to enable reproducibility and comparability across studies through benchmarking. High-frequency sampling should also be prioritized to produce longitudinal datasets that span entire growth cycles, improving the ability to model and predict temporal dynamics accurately. Furthermore, the development of multi-modal datasets, combining 2D images, 3D point cloud data, and environmental sensor measurements is suggested for capturing a holistic view of plant-environment interactions. Finally, integrating domain knowledge into dataset design to ensure that datasets not only capture observable traits but also reflect underlying biological processes. Such biologically informed datasets will enable models to generate predictions that are both empirically and biologically meaningful.

### v3

@@ 4.1 Functional-Structural Plant Models
Functional-structural plant models (FSPM) represent a class of plant representation models that construct the structural and geometric elements of a plant from its mechanistic processes [105, 106]. Particularly, these models capture the interaction between the plant’s physiological processes, such as photosynthesis, water transport, and nutrient allocation, and its development of plant organs such as leaves, branches, and roots over time [107, 108, 109, 110]. By coupling these two aspects, FSPMs provide a comprehensive view of plant behavior in response to various dependent parameters, allowing for the visualization of resource allocation and its effect on growth patterns. More recently, empirical methods have been integrated to the development process of FSPMs to establish the relationships between features extracted from HTPP data [111]. As such, an emergence of hybrid data-driven, mechanistic FSPMs have been applied towards plant growth simulations.
@@ 4.1.1 Process-driven Differential Growth Models
One application of these FSPMs is to capture the morphological developments within plant organs through differential growth models. The development and emergence of complex 3D morphologies within plant structures through time can be represented through numerical models, such as finite element methods (FEM), to describe differential growth across localized regions. The basis of the mechanistic models that determine these differentials are typically derived through experimental observations, utilizing image-processing techniques such as optical flow to obtain growth parameters from HTPP data [112]. Changes are mapped spatially and reconstructed from the distribution of conditions which influence the final shape of plant organs. One example of such modeling is demonstrated in Huang et al. [113], where biomechanical conditions are observed to determine the shape formations of leaves and petals. In this work, the modeling of growth differentials can be used to describe and reconstruct twisting, helical twisting, and edge-waving configurations within leaf structures through FEM simulations.
More often, this approach for plant structural modeling is applied toward developing fate and growth maps of cells within individual plant organs to simulate localized development [114]. For instance, in the work done by Kierzkowski et al. [115], the dynamics of genomic-proteomics pathways of two homeobox genes and their effects on promoting, inhibiting, and differentiating plant cells are modeled through spatial growth differentials using FEM. This phenomena is modeled through time and driven by the localized stimuli of plant hormone, namely auxin, concentrations. In this case, this model is applied toward developing a deep understanding of the causal effects of these specific signal pathways through empirical methods by mapping signals through time.
4.1.2 Rule-based Lindenmayer systems
Another common implementation of FSPMs involves the application of a rule-based modeling paradigm known as the Lindenmayer systems (L-systems). L-systems define a formal set of recursive rules that specify the configuration of plant structures based on a string of symbolic characters. These rules guide the growth and branching patterns of simulated plants through iterative processes that aim to generalize and represent the realistic biological development and propagation of plant structures [116, 117, 118]. By repeatedly applying these recursive rules, L-systems can simulate natural variations and intricate details found in plant morphology, making them a powerful tool in computational botany and generative modeling. A basic example is given by Algorithm 1, which establishes a generalized execution of a simple L-system, where Figure 6 shows the visual rendering of a rule set commonly used for describing a simple branching pattern in 2D.
@@ 4.1.2 Rule-based Lindenmayer systems
This algorithm can be expanded to include parameterized production rules that encapsulate complex relationships at each iteration of the simulated plant branching. For example, various works have developed production rules that dictate tree organ formation based on a source-sink dynamic between existing plant organs at previous iterations, which is then subsequently rendered for the 3D simulation of various tree cultivars [119, 120]. More recent work has expanded on production rules using sink-source dynamics by incorporating context-driven systems such as environmental and management conditions, soil nutrient availability, and distribution to improve the dynamic tree simulation process [121, 122, 123].
Furthermore, works such as Hitti et al. [124] and Mußmann et al. [125] introduce an adaptation of the L-systems paradigm that captures the dynamic development of plant structures using agent-based interactions. Plant structures, represented as agents, can independently interact with various stimuli. Within such systems, agent-based feedback mechanisms, such as reinforcement learning, can be used to train model growth patterns based on external stimuli.
@@ 4.2 Computer Vision Approach to Plant Modeling
FSPMs typically rely on procedural methods for the visual rendering process of the plant simulation pipeline (Figure 5). Although the outcomes of such models are interpretable and result in visually consistent structures, the outputs of such methods are often derived from deterministic solutions that lack stochastic variability. Such solutions tend to drift from biological realism, particularly when simulating complex interactions or environmental responses, requiring the integration of experiment data to drive realistic outputs [106]. Alternatively, some initial efforts have been made towards a novel approach using data-driven CV for rendering the structural elements of the simulation task.
@@ 4.2.1 Next Frame Prediction
Some early works on CV-based plant growth forecasting include Sakurai et al. [126] and Wang et al. [127], which have proposed the adoption of convolutional recurrent neural networks (CRNN) for the task of image-to-image translation to predict the next frame of a plant growth scene given a set of prior frames from the same scene. This approach expands on the sequence-based prediction techniques described in Section 3.2.2 to assess spatial features from consecutive images. This is done by incorporating convolutional layers within a RNN framework to extract and encode features from a set of input images, $X$, to a transitional hidden state $h$. The rendering of this hidden state is described by a decoding step which consists of an up-sampling process of $h$ through deconvolutional layers. The typical generic implementation of the CRNN is shown in Figure 7. This approach combines pattern recognition techniques for forecasting plant parameters using RNNs with the visual rendering of plant structures through image translation methods. While CRNNs allow for spatial features to be captured through time, the deterministic mapping through convolutional and deconvolutional layers often result in artifacts and ambiguities within structures of the output images [128].
To mitigate problems in image quality, recent work has focused on improving the image generation component of model architectures, which correspond to the decoder of a CRNN. Yasrab et al. [129] proposes the use of a Generative Adversarial Network (GAN)-based network architectures for improved realistic image synthesis. Other works such as Kim et al. [130] propose multi-step networks that separate various components of the image generation task, such as shape estimation and image reconstruction, utilizing a transformer-based network for spatial pattern recognition. Although these approaches improve image generation quality, they do not address the challenges of ambiguous structures caused by a deterministic mappings. CV-based plant forecasting is an ill-posed problem, in which, there exists non-unique outcomes of a plants growth trajectory given the same initial set of frames. This is particularly evident under diverging environmental conditions. Moreover, predictions from the current models are made solely on the previous frames of the scene and further contribute to the prevalence of ambiguities during image translation.
@@ 4.2.2 Frame Synthesis Through Conditional Generative Models
A probabilistic approach can be taken to regularize the plant growth modeling problem and to guide the prediction process by incorporating stochastic elements caused by dynamic growing conditions. In such methods, a generative model is employed for the task of frame synthesis to predict the expected frames of the plant growth scene. While generative models are typically used as data augmentation techniques for enhancing sparse agricultural datasets [131, 132], few works have explored the use of conditional generative models (CGM) for guided image synthesis. A probabilistic approach using a CGM is illustrated by a generic conditional variational autoencoder (CVAE) shown in figure 8. Instead of performing direct deterministic mappings between inputs and output frames, the objective of the CVAE is to learn a prior distribution, $p_{\theta}(z|x)$, where $z\sim p_{\theta}(z|x)$ is the latent representation of scene at a given point in time, $t$, directed by some conditional vector $x$. The prior distribution can be sampled such that it incorporates stochasticity through auxiliary variable, $\epsilon$. This process is known as latent sampling and it guides the generation process by randomly sampling from an underlying latent distribution that captures the hidden structure of the data. The network structures explored within the domain of plant growth modeling typically consists of a probabilistic latent variable encoder and a deterministic decoder to render output frames, $\hat{y}\sim p_{\theta}(y|z,x)$.
Early works, including Miranda et al. [133] and Drees et al. [134], propose the adoption of a conditional GAN (cGAN), which introduces a discriminator with adversarial routines to the generic CVAE structure, for synthesizing field images and forecasting the state of the crop canopy. The conditional vector consists of fixed soil treatment embeddings set at the start of the growing season. Forecasts are made at discrete points in time driven by the initial states of the treatment conditions. These works have demonstrated SOTA guided image generation for realistic image synthesis with variability in outputs defined by stochastic elements, or “factors of variation” (FOV). The frames synthesized from such models can be used in conjunction with image-based feature extraction models to provide quantitative estimation of biomass for a complete simulation process [135].
In these works, predictions are guided by a conditional vector comprised of treatment conditions, a prior frame state, and a time factor. The treatment embeddings are fixed at a particular point in time, typically set at the start of experimentation, and do not account for dynamic changes in conditions. For spatiotemporal simulation and trajectory estimation, the guidance of predictions without feedback can lead to larger variability when diverging from initial conditions due to the accumulated effects of compounding factors, including environmental conditions [136]. As such, this approach may be applicable for simulating images of end of harvest canopies under varying treatment conditions at a discrete point in time in the future, but lacks the dynamics for real-time continuous plant growth predictions useful for real world applications such as anomaly detection and intervention practices.
Debbagh et al. [137] proposes a sequence-informed approach which introduces dynamic and compounding environmental conditions as feedback to a CGM, guiding a continuous simulation process. In this work, the conditional vector, $x$, which guides the image generation of the CGM, is directed by a configuration of random variables, $c$, that describe dynamic factors that contribute to the development of the plants within the scene (i.e. environmental parameters). $x$ is determined by encoding a sequence of these environmental parameters from the start of the simulation up to the point of inference, $c(\tau)$, where $\tau=[0,t]$. The encoding process consists of an RNN encoder, which allows the model to learn the compounding and sequential relationships between environmental conditions as a vector representation. The dependencies between each model component is illustrated by the directed graph model shown in Figure 9. This approach provides a framework that further regularizes the image generation task through dynamic context, allowing for the realistic synthesis of consecutive frames.
4.3 Comparisons of Structured Output Representations Through Time
Table 4 provides a comparative summary of the structured modeling paradigms discussed in this section for simulating plant morphologies through time. Each method represents a trade-off between data needs, realism, and interpretability.
Mechanistic FSPMs such as differential growth simulations based on FEM rely on high-resolution imaging and cell-level observations to reconstruct localized growth processes. These models are highly interpretable, driven by physical and biological insights, but are typically deterministic and do not formally account for uncertainty. L-systems offer interpretable and symbolic modeling of branching structures using production rules. While L-systems are not probabilistic by default, stochastic extensions and agent-based variants allow for the incorporation of environmental interaction and feedback, enabling more dynamic simulations.
CV-based approaches introduce data-driven alternatives for visualizing plant growth. CRNNs are used to capture sequential spatial features from image sequences but suffer from low interpretability and deterministic output mappings, while CGMs, including the sequence-informed CVAE and cGANs, address these challenges by introducing latent variable modeling and sampling to represent uncertainty and generate more biologically plausible outputs.
Table 4: Comparison of Structured Models for Simulating Plant Morphologies Through Time
Model Data Requirements Uncertainty Handling Interpretability Application
Differential Growth (FEM) High; cell-level growth data, imaging-based flow analysis Implicit in parameter variability; deterministic model High; spatial and physical dynamics interpretable Leaf morphogenesis, biomechanical growth simulations
L-System Low to moderate; rule sets and symbolic encodings Not explicitly handled High; rules directly map to morphology Branching patterns, 2D/3D growth simulations, generative modeling
Agent-based L-System High; interaction data with agents/environment Stochasticity and feedback via agent logic; tunable Moderate to high; emergent from explicit interactions Context-aware growth modeling; reinforcement learning driven growth
CRNN (Next Frame Prediction) High; sequential image data of plant growth Implicit via temporal encoding Low to moderate; convolution and hidden state opacity Visual trait prediction; spatiotemporal structure forecasting
Conditional Generative Models (CVAE, cGAN) High; image pairs with metadata or environmental labels Explicit; latent distribution sampling (KL divergence, adversarial) Moderate; constrained latent space enables partial interpretation Visual scene generation; growth under static treatment conditions
  • Not probabilistic by default, but uncertainty can be modeled using Bayesian or variational extensions.
@@ 5.1 Data-Driven vs Experiment-Driven Approaches
Plant growth modeling has traditionally relied on experiment-driven approaches, where predefined hypotheses are tested using carefully controlled experiments. These methods provide valuable insights into specific biological processes but are often time-consuming, resource-intensive, and limited in scalability [138, 139, 140]. By contrast, data-driven and model-centric approaches leverage large datasets and machine learning techniques to capture patterns and relationships that might not be immediately apparent within an experiment-driven framework [141]. The works presented in this review demonstrate a shift towards the latter, as challenges with dynamic updating and long-term predictions are prevalent in simple regression methods as described in Section 3.1.3.
The shift towards 2D image and 3D point cloud data for describing plant phenemena, as with HTPP datasets, requires model frameworks designed for interpreting and processing higher order data-structures [142, 143]. This inherently moves prospective research towards the CV-based models described in Sections 4.2.1 and 4.2.2. However, the advantages of these data-driven approaches come with the limitation of reduced interpretability and a lack of the biological grounding inherent in experiment-driven methods [144]. Model frameworks such as FSPMs present potential opportunities for incorporating knowledge-based approaches for structured representation that is highly controllable, but is completely reliant on experiment-driven approaches for developing knowledge-based rule sets.
Future research in this domain should prioritize the development of hybrid approaches that integrate the strengths of both paradigms. Two key directions are proposed to achieve this goal. The first involves designing model architectures that incorporate knowledge-informed mechanisms, enabling the integration of biological principles and domain expertise into data-driven frameworks. This is demonstrated by works that aim to integrate deep learning frameworks into large scale crop models such as CERES and DSSAT [104]. The second focuses on advancing methods for decoding and interpreting opaque models, ensuring that complex representation models provide outputs that align with biological processes and facilitate deeper understanding [144].
@@ 5.2 Dataset Gap
The shift toward data-driven approaches has placed datasets at the center of plant growth modeling and predictive frameworks. Large-scale, high-quality datasets are necessary for training DL models and validating their predictions [145, 146]. However, the effectiveness of these models are dependent on the availability, diversity, and representativeness of the data [147]. Despite advancements in HTPP platforms, significant gaps remain in the scope, quality, and accessibility of existing datasets.
First, many datasets focus on single time points or static snapshots, which fail to capture the dynamic growth trajectories of plants over time. In such cases, works such as Zhu et al. [148] aim to simulate and analyze the physiological development of soybean plants using interpolated 3D reconstructions of discrete snapshots. Existing publicly available datasets that capture longitudinal data, compiled together and shown in Table 1, are scarce and limited in sample frequency. This is particularly evident for datasets that capture multiple modalities of data including 2D images and 3D point clouds along with environmental parameters. Second, the diversity of crop species represented by current datasets in Table 1 and HTTP platforms is limited, with a heavy emphasis on staple crops such as maize (Zea mays) and soybean (Glycine max) [149]. This narrow scope restricts the ability of models to generalize across plant species with differing morphologies, such as root vegetables or perennials. Finally, the lack of consolidation of datasets across different research groups and platforms creates significant barriers to broader adoption and cross-comparative studies. Inconsistent formats, annotation practices, and metadata standards hinder the ability to combine datasets, thereby limiting the diversity and scale necessary for developing generalizable models [150].
To address these gaps, several considerations for developing future datasets are proposed. First, works should aim to apply standardization and open-access practices during dataset development to enable reproducibility and comparability across studies through benchmarking. High-frequency sampling should also be prioritized to produce longitudinal datasets that span entire growth cycles, improving the ability to model and predict temporal dynamics accurately. Furthermore, the development of multi-modal datasets, combining 2D images, 3D point cloud data, and environmental sensor measurements is suggested for capturing a holistic view of plant-environment interactions. Finally, integrating domain knowledge into dataset design to ensure that datasets not only capture observable traits but also reflect underlying biological processes. Such biologically informed datasets will enable models to generate predictions that are both empirically and biologically meaningful [151].